OpenStudy (anonymous):

WILL GIVE MEDAL TO CORRECT ANSWER y^5 * y^14 =? 1. 2y^19 2.y^60 3. y^19 4. 2y^60

OpenStudy (anonymous):

3. y^19

OpenStudy (kainui):

So if you have x^3 what does that really mean? x*x*x. Now suppose you have x^3*x^2 what does that mean? (x^3)*(x^2)=(x*x*x)*(x*x)=(x*x*x*x*x)=x^5 right? So now you can see you just discovered adding exponents is the rule, simple as that!

OpenStudy (anonymous):

@Kainui So, it would be y^19?

OpenStudy (kainui):

Yep.

OpenStudy (anonymous):

Thanks!